Circular No. 6599

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

COMET C/1995 O1 (HALE-BOPP) J. E. Wink, Institut de Radio Astronomie Millimétrique (IRAM), Gre-

noble; D. Bockelée-Morvan, N. Biver, P. Colom, J. Crovisier, E. Gérard and H. Rauer, Observatoire de Paris; D. Despois, Observatoire de Bordeaux; R. Moreno and G. Paubert, IRAM, Granada; and J. K. Davies and W. R. F. Dent, Joint Astronomy Centre, Hilo, report the detection of formic acid at the IRAM Plateau de Bure interferometer using its five antennae in a single-dish mode: “On Mar. 20.6 UT the HCOOH rotational transitions at 224.977 GHz, 10(5,6)–9(5,5) and 10(5,5)–9(5,4), were detected with a total integrated line area on a T_{A}^{∗} scale of 0.08 ± 0.01 K km/s. The lines at 224.929 and 225.238 GHz were also detected. Assuming thermal excitation at 80 K, the derived HCOOH production rate is ∼ 5 × 10^{27} molecules/s, which is 50 times smaller than that of the related species CH_{3}OH.”

A. J. Apponi, T. C. Pesch and L. M. Ziurys, Steward Observatory, University of Arizona; and P. Wehinger and S. Wyckoff, Arizona State University, report the first detection of methyl cyanide in this comet: “On Mar. 20 the K = 0, 1, 2, 3 and 6 components of the J = 8–7 transition of CH_{3}CN near 147.1 GHz were observed using the NRAO 12-m telescope, with main beam brightness temperatures of 0.083, 0.059, 0.060, 0.081 and 0.053 K, respectively. Integrated intensities for the two strongest K components, K = 0 and K = 3, were 0.18 and 0.13 K km/s. In a search for vinyl cyanide (H_{2}CCHCN) at 151.9 GHz and ethyl cyanide (H_{3}CCH_{2}CN) at 152.3 GHz, no emission was detected to limits of 0.03 K peak-to-peak of main beam brightness temperature.”

GRS 1737−31

R. Sunyaev, E. Churazov, M. Revnivtsev and S. Trudolyubov, Space Research Institute, Moscow, and Max Planck Institute for Astrophysics, Garching, report for M. Vargas, J. Paul, J.-P. Roques, E. Jourdain and the GRANAT/SIGMA Team: “Since Mar. 14 GRANAT has been observing the Galactic Center region. Preliminary analysis of SIGMA hard x-ray images revealed a new source in Scorpius at α = 17^{h}36^{m}.7, δ = −30^{o}57′ (equinox 1950.0, error ±3′), 2^{o}.4 from the Galactic Center, at flux levels of 90±16 and 130±20 mCrab in the 35–75-keV and 75–150-keV bands, respectively. The photon index of the spectrum in this energy range is 1.6 ± 0.3. Assuming a distance of 8.5 kpc for the source, the hard x-ray luminosity is ∼ 10^{37} erg/s. The source was detected at practically the same flux level on Mar. 14.23– 15.59, 18.68–19.65 and 19.79–20.69 UT. The spectrum hardness suggests that this source is a distant x-ray nova and black-hole candidate.”

1997 March 21  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Brian G. Marsden